OpenStudy (anonymous):

Anyone can help ? Question 12.12. Which is the equation of the power function p(x) = ax3 where p(2) = 48? (Points : 4) p(x) = 24x3 p(x) = 6x3 p(x) = 6x3 p(x) = 8x3 Question 13.13. Which is the completely factored form of 27x3 + y6? (Points : 4) (3x + y2)3 (3x + y2)(9x2 + y4) (3x y2)(9x2 + 3xy2 + y4) (3x + y2)(9x2 3xy2 + y4) Question 14.14. Which function is an even function? (Points : 4) f(x) = x2 f(x) = 6x3 f(x) = x 2 f(x) = 2x3 + 2

OpenStudy (anonymous):

C o m e h e l p m e!!!!!!!!!!

OpenStudy (anonymous):

what is it involving ?

OpenStudy (anonymous):

Come to my question. scroll over my name and click what im veiwing.

OpenStudy (anonymous):

12.12 we look at the equation y(x)= ax^3; y(2) means we put to into all x's; y(2) = 48, therefore 48= a(2)^3, or a= 48/8, a=6. that means y(x)= 6x^2

OpenStudy (anonymous):

6x^3*

OpenStudy (anonymous):

thanks chris

OpenStudy (anonymous):

13.13 is D, it's hard to explain why, but if the space is a minus sign, the two equations are equal.

OpenStudy (anonymous):

14.14 is a very important question if your every going to continue in math, an even function means that regardless of the input being positive or negative, you will get the same answer, cos(x) is an even function, but sin(x) is not. This is because cos (60) and cos (-60) are both equal to 0.5. X^2 is an even function. Both 2^2 and (-2)^2 are 4. making x^2 an even function.

OpenStudy (anonymous):

are you sure its not -6x^3

OpenStudy (anonymous):

so the answer would be a ?

OpenStudy (anonymous):

I'm pretty sure that everything i wrote is right. 48/8 = 6, so for the first one y(x)=6x^3; (3x+y^2)(9x^2-3xy^2+y^4) would give you 27x^6 -9x^2y^2+9x^2y^2+3xy^4-3xy^4+y^6; and then the last one would be x^2 so A.